# A Survey on Large Language Models for Code Generation

### v1

@@ Abstract.
Large Language Models (LLMs) have garnered remarkable advancements across diverse code-related tasks, known as Code LLMs, particularly in code generation that generates source code with LLM from natural language descriptions. This burgeoning field has captured significant interest from both academic researchers and industry professionals due to its practical significance in software development, e.g., GitHub Copilot. Despite the active exploration of LLMs for a variety of code tasks, either from the perspective of natural language processing (NLP) or software engineering (SE) or both, there is a noticeable absence of a comprehensive and up-to-date literature review dedicated to LLM for code generation. In this survey, we aim to bridge this gap by providing a systematic literature review that serves as a valuable reference for researchers investigating the cutting-edge progress in LLMs for code generation. We introduce a taxonomy to categorize and discuss the recent developments in LLMs for code generation, covering aspects such as data curation, latest advances, performance evaluation, and real-world applications. In addition, we present a historical overview of the evolution of LLMs for code generation and offer an empirical comparison using the widely recognized HumanEval and MBPP benchmarks to highlight the progressive enhancements in LLM capabilities for code generation. We identify critical challenges and promising opportunities regarding the gap between academia and practical development. Furthermore, we have established a dedicated resource website (https://codellm.github.io) to continuously document and disseminate the most recent advances in the field.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 9ccs: General and reference Surveys and overviewsccs: Software and its engineering Software development techniquesccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
The advent of Large Language Models (LLMs) such as ChatGPT [^1] https://chat.openai.com (OpenAI, 2022) has profoundly transformed the landscape of automated code-related tasks (Chen et al., 2021), including code completion (Wang and Li, 2021; Lu et al., 2022; Guo et al., 2023; Wu et al., 2024), code translation (Lachaux et al., 2020; Szafraniec et al., 2022; Chen et al., 2018), and code repair (Jiang et al., 2021; Olausson et al., 2023; Parasaram et al., 2024). A particularly intriguing application of LLMs is code generation, a task that involves producing source code from natural language descriptions. Despite varying definitions across studies (Ren et al., 2020; Chen et al., 2023; Shojaee et al., 2023; Wang et al., 2023b), for the purposes of this survey, we adopt a consistent definition of code generation as the natural-language-to-code (NL2Code) task (Austin et al., 2021; Athiwaratkun et al., 2022; Zan et al., 2023). This area has garnered substantial interest from both academia and industry, as evidenced by the development of tools like GitHub Copilot [^2] https://github.com/features/copilot (Chen et al., 2021), CodeGeeX [^3] https://codegeex.cn/en-US (Zheng et al., 2023c), and Amazon CodeWhisperer [^4] https://aws.amazon.com/codewhisperer, which leverage groundbreaking code LLMs to facilitate software development.
Initial investigations into code generation primarily utilized heuristic rules or expert systems, such as probabilistic grammar-based frameworks (Joshi and Rambow, 2003; Cohn et al., 2010; Allamanis and Sutton, 2014) and specialized language models (De Moura and Bjørner, 2008; Gulwani, 2010; Jha et al., 2010). These early techniques were typically rigid and difficult to scale. However, the introduction of Transformer-based LLMs has shifted the paradigm, establishing them as the preferred method due to their superior proficiency and versatility. One remarkable aspect of LLMs is their capability to follow instructions (Wei et al., 2022a; Ouyang et al., 2022; Xu et al., 2023; Muennighoff et al., 2023; Chung et al., 2024), enabling even novice programmers to write code by simply articulating their requirements. This emergent ability has democratized coding, making it accessible to a broader audience (Zan et al., 2023). The performance of LLMs on code generation tasks has seen remarkable improvements, as illustrated by the HumanEval leaderboard [^5] https://paperswithcode.com/sota/code-generation-on-humaneval, which showcases the evolution from PaLM 8B (Chowdhery et al., 2023) of 3.6% to LDB (Zhong et al., 2024) of 95.1% on Pass@1 metrics. As can be seen, the HumanEval benchmark (Chen et al., 2021) has been established as a de facto standard for evaluating the coding proficiency of LLMs (Chen et al., 2021).
@@
While recent surveys have shed light on code LLMs from the lenses of Natural Language Processing (NLP), Software Engineering (SE), or a combination of both disciplines (Zan et al., 2023; Zheng et al., 2023a; Zhang et al., 2023b; Hou et al., 2024), they have often encompassed a broad range of code-related tasks. There remains a dearth of literature specifically reviewing advanced topics in code generation, such as meticulous data curation, instruction tuning, alignment with feedback, prompting techniques, the development of autonomous coding agents, retrieval augmented code generation, LLM-as-a-Judge for code generation, among others. A notably pertinent study (Athiwaratkun et al., 2022; Zan et al., 2023) also concentrates on LLMs for text-to-code generation (NL2Code), yet it primarily examines models released from 2020 to 2022. Consequently, this noticeable temporal gap has resulted in an absence of up-to-date literature reviews that contemplate the latest advancements, including models like CodeQwen (Team, 2024), WizardCoder (Luo et al., 2023), and PPOCoder (Shojaee et al., 2023), as well as the comprehensive exploration of the advanced topics previously mentioned.
@@
Recognizing the need for a dedicated and up-to-date literature review, this survey endeavors to fill that void. We provide a systematic review that will serve as a foundational reference for researchers quickly exploring the latest progress in LLMs for code generation. A taxonomy is introduced to categorize and examine recent advancements, encompassing data curation (Wang et al., 2023a; Luo et al., 2023; Wei et al., 2023), advanced topics (Parvez et al., 2021; Lu et al., 2022; Le et al., 2022; Muennighoff et al., 2023; Liu et al., 2023d; Chen et al., 2022b; Ni et al., 2023a; Chen et al., 2023; Huang et al., 2023; Shrivastava et al., 2023a; Zhang et al., 2023c), evaluation methods (Chen et al., 2021; Hendrycks et al., 2021; Jimenez et al., 2023; Zhuo, 2024), and practical applications (Chen et al., 2021; Zheng et al., 2023c). This category aligns with the comprehensive lifecycle of an LLM for code generation. Furthermore, we pinpoint critical challenges and identify promising opportunities to bridge the research-practicality divide. Therefore, this survey allows NLP and SE researchers to seamlessly equip with a thorough understanding of LLM for code generation, highlighting cutting-edge directions and current hurdles and prospects.
@@
The remainder of the survey is organized following the structure outlined in our taxonomy in Figure 3. In Section 2, we introduce the preliminaries of LLM with Transformer architecture and formulate the task of LLM for code generation. Then, in Section 3, we propose a taxonomy, categorizing the complete process of LLMs in code generation. Section 4 delves into the specifics of LLMs for code generation within this taxonomy framework. In Section 5, we underscore the critical challenges and promising opportunities for bridging the research-practicality gap and conclude this work in Section 6.
@@ 2.1. Large Language Models
The effectiveness of large language models (LLMs) is fundamentally attributed to their substantial quantity of model parameters, large-scale and diversified datasets, and the immense computational power utilized during training (Kaplan et al., 2020; Hoffmann et al., 2022). Generally, scaling up language models consistently results in enhanced performance and sample efficiency across a broad array of downstream tasks (Wei et al., 2022a; Zhao et al., 2023b). However, with the expansion of the model size to a certain extent (e.g., GPT-3 (Brown et al., 2020) with 175B-parameters and PaLM (Chowdhery et al., 2023) with 540B), LLMs have exhibited an unpredictable phenomenon known as emergent abilities [^6] It should be noted that an LLM is not necessarily superior to a smaller language model, and emergent abilities may not manifest in all LLMs (Zhao et al., 2023b)., including instruction following (Ouyang et al., 2022), in-context learning (Dong et al., 2022), and step-by-step reasoning (Wei et al., 2022b; Huang and Chang, 2022), which are absent in smaller models but apparent in larger ones (Wei et al., 2022a).
Adhering to the same architectures of the Transformer (Vaswani et al., 2017) in LLMs, code LLMs are specifically pre-trained on large-scale unlabeled code corpora, whereas general-purpose LLMs (e.g., ChatGPT (OpenAI, 2022)) are pre-trained on a blend of code and text data. Analogous to LLMs, Code LLMs can also be classified into three architectural categories: encoder-only models, decoder-only models, and encoder-decoder models. For encoder-only models, such as CodeBERT (Feng et al., 2020), they are typically suitable for code comprehension tasks including type prediction, code retrieval, and clone detection. For decoder-only models, such as StarCoder (Brown et al., 2020), they predominantly excel in generation tasks, such as code generation, code translation, and code summarization. Encoder-decoder models, such as CodeT5 (Wang et al., 2021), can accommodate both code understanding and generation tasks but do not necessarily outperform encoder-only or decoder-only models. The overall architectures of the different Code LLMs for code generation are depicted in Figure 2.
@@ 2.1.1. Multi-Head Self-Attention Modules
where $\mathbf{H}^{(l-1)}\in\mathbb{R}^{n\times d_{model}}$ denotes the input to the $l$-th Transformer layer, while $\mathbf{h}^{(l)}\in\mathbb{R}^{n\times d_{model}}$ represents the output of MHSA sub-layer. The quantity of distinct attention heads is represented by $h$, and $d_{model}$ refers to the model dimension. The set of projections $\left\{\mathbf{W}_{i}^{\mathbf{Q}},\mathbf{W}_{i}^{\mathbf{K}},\mathbf{W}_{i}^{\mathbf{V}},\mathbf{W}_{i}^{\mathbf{O}}\right\}\in\mathbb{R}^{d_{model}\times d_{model}/h}$ encompasses the affine transformation parameters for each attention head $\operatorname{Head}_{i}$, transforming the Query $\mathbf{Q}$, Key $\mathbf{K}$, Value $\mathbf{V}$, and the output of the attention sub-layer, The $\operatorname{softmax}$ function is applied in a row-wise manner. The dot-products of queries and keys are divided by a scaling factor $\sqrt{d_{model}/h}$ to counteract the potential risk of excessive large inner products and correspondingly diminished gradients in the $\operatorname{softmax}$ function, thus encouraging a more balanced attention landscape.
@@ 2.2. Code Generation
Large language models (LLMs) for code generation refer to the use of LLM to generate source code from natural language descriptions, a process also known as a natural-language-to-code task. Typically, these natural language descriptions encompass programming problem statements (or docstrings) and may optionally include some programming context (e.g., function signatures, assertions, etc.). Formally, these natural language (NL) descriptions can be represented as $\mathbf{x}$. Given $\mathbf{x}$, the use of an LLM with model parameters $\theta$ to generate a code solution $\mathbf{y}$ can be denoted as $P_{\theta}(\mathbf{y}\mid\mathbf{x})$. To verify the functionality correctness of the code solution, $\mathbf{y}$ is subsequently executed via a compiler or interpreter, represented as $\mathbf{Exe}(\cdot)$, on a suit of unit tests. The feedback from this execution can be denoted as $\mathbf{Feedback}(\mathbf{Exe}(\mathbf{y}))$.
The advent of in-context learning abilities in LLM (Wei et al., 2022a) has led to the appending of exemplars to the natural language description $\mathbf{x}$ as demonstrations to enhance code generation performance or constrain the generation format (Li et al., 2023d; Patel et al., 2023). A fixed set of $M$ exemplars is denoted as $\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{M}$. Consequently, following (Ni et al., 2023a), a more general formulation of LLMs for code generation with few-shot (or zero-shot) exemplars can be revised as:
(12) $P_{\theta}(\mathbf{y}\mid\mathbf{x})=P_{\theta}(\mathbf{y}\mid\operatorname{prompt}(\mathbf{x},\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{k})),k=\{0,1,\dots,M\}$
@@
where $\operatorname{prompt}(\mathbf{x},\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{k}))$ is a string representation of the overall input, and $\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{k}$ denotes a set of $k$ exemplars randomly selected from $\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{M}$. In particular, when $k=0$, this denotes zero-shot code generation, equivalent to vanilla ones without in-context learning. Subsequently, a variety of decoding strategies can be performed for code generation, including deterministic-based strategies (e.g., greedy search and beam search) and sampling-based strategies (e.g., temperature sampling, top-k sampling, and top-p (nucleus) sampling). For more detailed descriptions of each decoding strategy, please consult (Holtzman et al., 2019).

### v2

@@ Abstract.
Large Language Models (LLMs) have garnered remarkable advancements across diverse code-related tasks, known as Code LLMs, particularly in code generation that generates source code with LLM from natural language descriptions. This burgeoning field has captured significant interest from both academic researchers and industry professionals due to its practical significance in software development, e.g., GitHub Copilot. Despite the active exploration of LLMs for a variety of code tasks, either from the perspective of natural language processing (NLP) or software engineering (SE) or both, there is a noticeable absence of a comprehensive and up-to-date literature review dedicated to LLM for code generation. In this survey, we aim to bridge this gap by providing a systematic literature review that serves as a valuable reference for researchers investigating the cutting-edge progress in LLMs for code generation. We introduce a taxonomy to categorize and discuss the recent developments in LLMs for code generation, covering aspects such as data curation, latest advances, performance evaluation, ethical implications, environmental impact, and real-world applications. In addition, we present a historical overview of the evolution of LLMs for code generation and offer an empirical comparison using the HumanEval, MBPP, and BigCodeBench benchmarks across various levels of difficulty and types of programming tasks to highlight the progressive enhancements in LLM capabilities for code generation. We identify critical challenges and promising opportunities regarding the gap between academia and practical development. Furthermore, we have established a dedicated resource GitHub page (https://github.com/juyongjiang/CodeLLMSurvey) to continuously document and disseminate the most recent advances in the field.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 1publicationmonth: 8ccs: General and reference Surveys and overviewsccs: Software and its engineering Software development techniquesccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
The advent of Large Language Models (LLMs) such as ChatGPT [^1] https://chat.openai.com (OpenAI, 2022) has profoundly transformed the landscape of automated code-related tasks (Chen et al., 2021), including code completion (Wang and Li, 2021; Lu et al., 2022; Guo et al., 2023; Wu et al., 2024), code translation (Lachaux et al., 2020; Szafraniec et al., 2022; Chen et al., 2018), and code repair (Olausson et al., 2023; Fan et al., 2023a; Joshi et al., 2023; Parasaram et al., 2024; Xu et al., 2024a; Zhang et al., 2024a). A particularly intriguing application of LLMs is code generation, a task that involves producing source code from natural language descriptions. Despite varying definitions across studies (Ren et al., 2020; Chen et al., 2023; Shojaee et al., 2023; Wang et al., 2023b), for the main scope of this survey, we focus on the code generation task and adopt a consistent definition of code generation as the natural-language-to-code (NL2Code) task (Austin et al., 2021; Athiwaratkun et al., 2022; Zan et al., 2023). To enhance clarity, the differentiation between code generation and other code-related tasks, along with a more nuanced definition, is summarized in Table 1. This area has garnered substantial interest from both academia and industry, as evidenced by the development of tools like GitHub Copilot [^2] https://github.com/features/copilot (Chen et al., 2021), CodeGeeX [^3] https://codegeex.cn/en-US (Zheng et al., 2023c), and Amazon CodeWhisperer [^4] https://aws.amazon.com/codewhisperer, which leverage groundbreaking code LLMs to facilitate software development.
Initial investigations into code generation primarily utilized heuristic rules or expert systems, such as probabilistic grammar-based frameworks (Joshi and Rambow, 2003; Cohn et al., 2010; Allamanis and Sutton, 2014; Xiong et al., 2017; Ji et al., 2020) and specialized language models (De Moura and Bjørner, 2008; Gulwani, 2010; Jha et al., 2010). These early techniques were typically rigid and difficult to scale. However, the introduction of Transformer-based LLMs has shifted the paradigm, establishing them as the preferred method due to their superior proficiency and versatility. One remarkable aspect of LLMs is their capability to follow instructions (Wei et al., 2022a; Ouyang et al., 2022; Xu et al., 2023; Muennighoff et al., 2023; Chung et al., 2024), enabling even novice programmers to write code by simply articulating their requirements. This emergent ability has democratized coding, making it accessible to a broader audience (Zan et al., 2023). The performance of LLMs on code generation tasks has seen remarkable improvements, as illustrated by the HumanEval leaderboard [^5] https://paperswithcode.com/sota/code-generation-on-humaneval, which showcases the evolution from PaLM 8B (Chowdhery et al., 2023) of 3.6% to LDB (Zhong et al., 2024) of 95.1% on Pass@1 metrics. As can be seen, the HumanEval benchmark (Chen et al., 2021) has been established as a de facto standard for evaluating the coding proficiency of LLMs (Chen et al., 2021).
@@
While recent surveys have shed light on code LLMs from the lenses of Natural Language Processing (NLP), Software Engineering (SE), or a combination of both disciplines (Zan et al., 2023; Zheng et al., 2023a; Zhang et al., 2023b; Fan et al., 2023b; Hou et al., 2024; Lyu et al., 2024), they have often encompassed a broad range of code-related tasks. There remains a dearth of literature specifically reviewing advanced topics in code generation, such as meticulous data curation, instruction tuning, alignment with feedback, prompting techniques, the development of autonomous coding agents, retrieval augmented code generation, LLM-as-a-Judge for code generation, among others. A notably pertinent study (Athiwaratkun et al., 2022; Zan et al., 2023) also concentrates on LLMs for text-to-code generation (NL2Code), yet it primarily examines models released from 2020 to 2022. Consequently, this noticeable temporal gap has resulted in an absence of up-to-date literature reviews that contemplate the latest advancements, including models like CodeQwen (Team, 2024), WizardCoder (Luo et al., 2023), CodeFusion (Singh et al., 2023), and PPOCoder (Shojaee et al., 2023), as well as the comprehensive exploration of the advanced topics previously mentioned.
Table 1. The applications of code LLMs in various code-related understanding and generation tasks. The I-O column indicates the type of input and output for each task, where C, NL, and K represent code, natural language, and label, respectively. Note that the detailed definitions of each task aligns with the descriptions in (Ahmad et al., 2021; Austin et al., 2021; Niu et al., 2022; Athiwaratkun et al., 2022; Zan et al., 2023). The main scope of this survey focuses on code generation while it may involve code completion in Section 5.7 and 5.8, aiming to illustrate the corresponding advancements.
Type I-O Task Definition
Understanding C-K Code Classification Classify code snippets based on functionality, purpose, or attributes to aid in organization and analysis.
Bug Detection Detect and diagnose bugs or vulnerabilities in code to ensure functionality and security.
Clone Detection Identifying duplicate or similar code snippets in software to enhance maintainability, reduce redundancy, and check plagiarism.
Exception Type Prediction Predict different exception types in code to manage and handle exceptions effectively.
C-C Code-to-Code Retrieval Retrieve relevant code snippets based on a given code query for reuse or analysis.
NL-C Code Search Find relevant code snippets based on natural language queries to facilitate coding and development tasks.
Generation C-C Code Completion Predict and suggest the next portion of code, given contextual information from the prefix (and suffix), while typing to enhance development speed and accuracy.
Code Translation Translate the code from one programming language to another while preserving functionality and logic.
Code Repair Identify and fix bugs in code by generating the correct version to improve functionality and reliability.
Mutant Generation Generate modified versions of code to test and evaluate the effectiveness of testing strategies.
Test Generation Generate test cases to validate code functionality, performance, and robustness.
C-NL Code Summarization Generate concise textual descriptions or explanations of code to enhance understanding and documentation.
NL-C Code Generation Generate source code from natural language descriptions to streamline development and reduce manual coding efforts.
@@
Recognizing the need for a dedicated and up-to-date literature review, this survey endeavors to fill that void. We provide a systematic review that will serve as a foundational reference for researchers quickly exploring the latest progress in LLMs for code generation. A taxonomy is introduced to categorize and examine recent advancements, encompassing data curation (Wang et al., 2023a; Luo et al., 2023; Wei et al., 2023b), advanced topics (Parvez et al., 2021; Lu et al., 2022; Le et al., 2022; Muennighoff et al., 2023; Liu et al., 2023e; Chen et al., 2022c; Ni et al., 2023a; Chen et al., 2023; Huang et al., 2023; Shrivastava et al., 2023a; Zhang et al., 2023c), evaluation methods (Chen et al., 2021; Hendrycks et al., 2021; Jimenez et al., 2023; Zhuo, 2024), and practical applications (Chen et al., 2021; Zheng et al., 2023c). This category aligns with the comprehensive lifecycle of an LLM for code generation. Furthermore, we pinpoint critical challenges and identify promising opportunities to bridge the research-practicality divide. Therefore, this survey allows NLP and SE researchers to seamlessly equip with a thorough understanding of LLM for code generation, highlighting cutting-edge directions and current hurdles and prospects.
@@
The remainder of the survey is organized following the structure outlined in our taxonomy in Figure 6. In Section 2, we introduce the preliminaries of LLM with Transformer architecture and formulate the task of LLM for code generation. Section 3, we detail the systematic methodologies employed in conducting literature reviews. Then, in Section 4, we propose a taxonomy, categorizing the complete process of LLMs in code generation. Section 5 delves into the specifics of LLMs for code generation within this taxonomy framework. In Section 6, we underscore the critical challenges and promising opportunities for bridging the research-practicality gap and conclude this work in Section 7.
@@ 2.1. Large Language Models
The effectiveness of large language models (LLMs) is fundamentally attributed to their substantial quantity of model parameters, large-scale and diversified datasets, and the immense computational power utilized during training (Kaplan et al., 2020; Hoffmann et al., 2022). Generally, scaling up language models consistently results in enhanced performance and sample efficiency across a broad array of downstream tasks (Wei et al., 2022a; Zhao et al., 2023b). However, with the expansion of the model size to a certain extent (e.g., GPT-3 (Brown, 2020) with 175B-parameters and PaLM (Chowdhery et al., 2023) with 540B), LLMs have exhibited an unpredictable phenomenon known as emergent abilities [^6] It should be noted that an LLM is not necessarily superior to a smaller language model, and emergent abilities may not manifest in all LLMs (Zhao et al., 2023b)., including instruction following (Ouyang et al., 2022), in-context learning (Dong et al., 2022), and step-by-step reasoning (Wei et al., 2022b; Huang and Chang, 2022), which are absent in smaller models but apparent in larger ones (Wei et al., 2022a).
Adhering to the same architectures of the Transformer (Vaswani et al., 2017) in LLMs, code LLMs are specifically pre-trained (or continually pre-trained on general LLMs) using large-scale unlabeled code corpora with a smaller portion of text (and math) data, whereas general-purpose LLMs are pre-trained primarily on large-scale text data, incorporating a smaller amount of code and math data to enhance logical reasoning capabilities. Additionally, some code LLMs, such as Qwen2.5-Coder (Hui et al., 2024), incorporate synthetic data in their training processes, a practice that is attracting increasing attention from both industry and academia. Analogous to LLMs, Code LLMs can also be classified into three architectural categories: encoder-only models, decoder-only models, and encoder-decoder models. For encoder-only models, such as CodeBERT (Feng et al., 2020), they are typically suitable for code comprehension tasks including type prediction, code retrieval, and clone detection. For decoder-only models, such as StarCoder (Brown, 2020), they predominantly excel in generation tasks, such as code generation, code translation, and code summarization. Encoder-decoder models, such as CodeT5 (Wang et al., 2021), can accommodate both code understanding and generation tasks but do not necessarily outperform encoder-only or decoder-only models. The overall architectures of the different Code LLMs for code generation are depicted in Figure 2.
@@ 2.1.1. Multi-Head Self-Attention Modules
where $\mathbf{H}^{(l-1)}\in\mathbb{R}^{n\times d_{model}}$ denotes the input to the $l$-th Transformer layer, while $\mathbf{h}^{(l)}\in\mathbb{R}^{n\times d_{model}}$ represents the output of MHSA sub-layer. The quantity of distinct attention heads is represented by $h$, and $d_{model}$ refers to the model dimension. The set of projections $\left\{\mathbf{W}_{i}^{\mathbf{Q}},\mathbf{W}_{i}^{\mathbf{K}},\mathbf{W}_{i}^{\mathbf{V}},\mathbf{W}_{i}^{\mathbf{O}}\right\}\in\mathbb{R}^{d_{model}\times d_{model}/h}$ encompasses the affine transformation parameters for each attention head $\operatorname{Head}_{i}$, transforming the Query $\mathbf{Q}$, Key $\mathbf{K}$, Value $\mathbf{V}$, and the output of the attention sub-layer. The $\operatorname{softmax}$ function is applied in a row-wise manner. The dot-products of queries and keys are divided by a scaling factor $\sqrt{d_{model}/h}$ to counteract the potential risk of excessive large inner products and correspondingly diminished gradients in the $\operatorname{softmax}$ function, thus encouraging a more balanced attention landscape.
@@ 2.1.4. Positional Encoding
2.1.5. Architecture
There are two types of Transformer architecture for code generation task, including encoder-decoder and decoder-only. For the encoder-decoder architecture, it consists of both an encoder and a decoder, in which the encoder processes the input data and generates a set of representations, which are then used by the decoder to produce the output. However, for decoder-only architecture, it consists only of the decoder part of the transformer, where it uses a single stack of layers to both process input data and generate output. Therefore, the encoder-decoder architecture is suited for tasks requiring mapping between different input and output domains, while the decoder-only architecture is designed for tasks focused on sequence generation and continuation. The overview of LLMs with these two architectures are illustrated in Figure 2.
@@ 2.2. Code Generation
Large language models (LLMs) for code generation refer to the use of LLM to generate source code from natural language descriptions, a process also known as a natural-language-to-code task. Typically, these natural language descriptions encompass programming problem statements (or docstrings) and may optionally include some programming context (e.g., function signatures, assertions, etc.). Formally, these natural language (NL) descriptions can be represented as $\mathbf{x}$. Given $\mathbf{x}$, the use of an LLM with model parameters $\theta$ to generate a code solution $\mathbf{y}$ can be denoted as $P_{\theta}(\mathbf{y}\mid\mathbf{x})$. The advent of in-context learning abilities in LLM (Wei et al., 2022a) has led to the appending of exemplars to the natural language description $\mathbf{x}$ as demonstrations to enhance code generation performance or constrain the generation format (Li et al., 2023d; Patel et al., 2023). A fixed set of $M$ exemplars is denoted as $\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{M}$. Consequently, following (Ni et al., 2023a), a more general formulation of LLMs for code generation with few-shot (or zero-shot) exemplars can be revised as:
(12) $P_{\theta}(\mathbf{y}\mid\mathbf{x})\Rightarrow P_{\theta}(\mathbf{y}\mid\operatorname{prompt}(\mathbf{x},\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{k})),k\in\{0,1,\dots,M\}$
@@
where $\operatorname{prompt}(\mathbf{x},\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{k}))$ is a string representation of the overall input, and $\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{k}$ denotes a set of $k$ exemplars randomly selected from $\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{M}$. In particular, when $k=0$, this denotes zero-shot code generation, equivalent to vanilla ones without in-context learning. In the decoding process, a variety of decoding strategies can be performed for code generation, including deterministic-based strategies (e.g., greedy search and beam search) and sampling-based strategies (e.g., temperature sampling, top-k sampling, and top-p (nucleus) sampling). For more detailed descriptions of each decoding strategy, please consult (Holtzman et al., 2019). For example, the greedy search and sampling-based decoding strategies can be formulated as follows: